OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sin(5pi/12)

OpenStudy (anonymous):

@zzr0ck3r @SithsAndGiggles @surjithayer

OpenStudy (anonymous):

fore warning, i don't know what half angle identity is

OpenStudy (mathmate):

You'll find it here: http://home.windstream.net/okrebs/page103.html \(\huge sin(x)=\pm sqrt{\frac{(1-cos(2x))}{2}} \)

OpenStudy (mathmate):

Set x=5pi/12, then 2x=10pi/12=5pi/6, and cos(5pi/6)=cos(pi/6)=sqrt(3)/2 Take the positive value of sin(x) because sin(x)>0 for all 0<x<pi.

OpenStudy (anonymous):

\[\sin \left( \frac{ 5\pi }{ 12 } \right)=\sqrt{\frac{ 1-\cos \left( \frac{ 5\pi }{ 6 } \right)}{ 2 }}\] so it should look like this?

OpenStudy (mathmate):

Yes. Note that I made a mistake above. cos(5pi/6)=-pi/6=-sqrt(3)/2 So that will change slightly if you decide to replace cos(5pi/6) by cos(pi/6).

OpenStudy (anonymous):

\[\sin \left( \frac{ 5\pi }{ 12 } \right)=\sqrt{\frac{ 1-\cos \left( \frac{ \pi }{ 6 } \right) }{ 12 }}\]

OpenStudy (anonymous):

that second 12 should be a 2